# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v1

@@ Appendix A Online uncertainty updates: Kalman–Bucy filtering and a plug-in decision-time benchmark
Remarks (relation to belief-aware control). The plug-in rule (96) is intentionally decision-time: it conditions on the current uncertainty description and does not attempt to optimize over how the posterior will evolve. In the present paper, we therefore treat the uncertainty law as fixed at a decision time (either as a fixed $q$ over a horizon, or as an externally updated sequence of inputs $q_{t}$ that is not controlled by the agent), mirroring the decision-time perspective common in Bayesian/learning portfolio-choice studies (barberis2000investor; pastor2000portfolio; xia2001learning). Even in linear–Gaussian regimes where the belief state is finite-dimensional, the fully optimal partial-observation portfolio problem would treat the belief state (here, $(\hat{Y}_{t},P_{t})$) as part of the controlled state and optimize the policy in that enlarged state space (bensoussan1985optimal; pham2017dynamic). Related necessary conditions under partial information can also be expressed via partial-observation maximum principles (haussmann1987maximum; li1995general; baghery2007maximum). Developing a belief-aware Pontryagin-guided policy optimizer that operates directly in $(x,y,\hat{Y},P)$-space (or its sufficient-statistic analogues) is an important direction that we defer to future work.
Appendix B Proof of Theorem 2
@@
Theorem 2 extends the BPTT–PMP (equivalently, BPTT–BSDE) correspondence established for deterministic-parameter models in our prior work on PG–DPO (see the main BPTT–BSDE correspondence result and proof in huh2025breaking). Here the only substantive change is that the market coefficients are indexed by a random but frozen parameter $\theta\sim q$, and we need convergence statements that hold conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$.
@@
Important remark (what this proof does not use). This proof concerns the $\theta$-conditional Pontryagin adjoint/costate for the fixed-$\theta$ control problem induced by (53)–(54). It does not use the $\theta$-blind $q$-aggregated stationarity condition (Theorem 1 in Section 3.2). Those constructions affect only the deployable aggregation/projection target in stage 2 and are irrelevant to the BPTT–PMP convergence itself.
@@
Notation and filtration. Fix $\theta\in\Theta$. We work conditionally on this $\theta$ and consider the augmented (simulator) filtration
@@
$\mathbb{G}^{\theta}:=(\mathcal{G}_{t}^{\theta})_{t\in[0,T]},\hskip 18.49988pt\mathcal{G}_{t}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t\}\big)\ \text{(with the usual augmentation)}.$
All conditional expectations and $L^{2}$ projections below are taken with respect to $\mathcal{G}_{t_{k}}^{\theta}$. This choice matches the information set used by the simulator and by automatic differentiation/BPTT (which differentiates through the full forward recursion).
@@
Let $\Delta t>0$, $t_{k}:=k\Delta t$, $k=0,\dots,N$, $N\Delta t=T$. For readability we suppress the policy parameters $\varphi$ and write $\pi_{k}:=\pi_{\varphi}(t_{k},X_{k}^{\theta},Y_{k}^{\theta})$, where $\pi_{\varphi}$ is $\theta$-blind in the sense of (52) but evaluated along the $\theta$-conditional trajectory.
@@
Step 1: Conditioning on $\theta$ and uniformity of bounds. Fix a compact set $K\subset\Theta$. Assume the coefficients in (53)–(54) satisfy the usual Lipschitz and linear-growth conditions uniformly over $\theta\in K$, and that the block covariance structure of $(W,W^{Y})$ (including instantaneous correlation) is uniformly nondegenerate on $K$. Then, for each fixed $\theta\in K$, the controlled SDE system (53)–(54) is well posed and admits uniform-in-time $L^{2}$ moment bounds. Moreover, the Euler–Maruyama scheme enjoys the standard strong error bound
$\sup_{t\in[0,T]}\mathbb{E}\big[\|(X_{t}^{\pi,\theta},Y_{t}^{\theta})-(X_{t}^{\Delta t,\theta},Y_{t}^{\Delta t,\theta})\|^{2}\big]^{1/2}\;\leq\;C_{K}\,\Delta t^{1/2},$
@@
with a constant $C_{K}$ that can be chosen independently of $\theta\in K$. These are the deterministic assumptions used in huh2025breaking, now stated uniformly on $K$.
@@
Step 2: Discrete forward scheme and BPTT pathwise adjoints (fixed $\theta$). Under fixed $\theta$, consider the Euler scheme for (53)–(54) on the grid $(t_{k})$:
@@
with terminal reward $U(X_{N}^{\theta})$. Define the discrete (pathwise) wealth costate
@@
which is the same object as (59) (episode indices suppressed and dependence on $\theta$ made explicit). For the projected-control constructions we also consider the additional pathwise objects
$p_{x,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial X_{k}^{\theta}},\hskip 18.49988ptp_{y,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial Y_{k}^{\theta}},$
which correspond to (60). Automatic differentiation/BPTT computes $\{(p_{k}^{\mathrm{pw},\theta},p_{x,k}^{\mathrm{pw},\theta},p_{y,k}^{\mathrm{pw},\theta})\}_{k=0}^{N}$ via the backward chain rule along the discrete forward graph.
The algebraic form of the one-step backward recursion coincides with the deterministic-parameter analysis in huh2025breaking, with the replacements
@@
and with the factor block $(Y^{\theta},\Delta W^{Y})$ handled exactly as in the wealth–factor extension therein. All one-step remainder terms are controlled by standard Taylor/Euler estimates with constants uniform in $\theta\in K$.
Step 3: One-step $L^{2}$ projection and discrete BSDE form (fixed $\theta$). Fix $\theta\in K$. As in huh2025breaking, take the conditional $L^{2}$-projection of $p_{k+1}^{\mathrm{pw},\theta}$ onto $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ given $\mathcal{G}_{t_{k}}^{\theta}$:
@@
$p_{k+1}^{\mathrm{pw},\theta}=\mathbb{E}\!\left[p_{k+1}^{\mathrm{pw},\theta}\mid\mathcal{G}_{t_{k}}^{\theta}\right]+z_{k}^{\theta}\Delta W_{k}+\tilde{z}_{k}^{\theta}\Delta W_{k}^{Y}+\varepsilon_{k+1}^{\theta},$
where $\varepsilon_{k+1}^{\theta}$ is orthogonal (in $L^{2}$) to $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ conditionally on $\mathcal{G}_{t_{k}}^{\theta}$. Uniform nondegeneracy of the block covariance of $(\Delta W_{k},\Delta W_{k}^{Y})$ yields unique projection coefficients $(z_{k}^{\theta},\tilde{z}_{k}^{\theta})$.
@@
Substituting this projection into the BPTT backward recursion from Step 2 yields a canonical discrete BSDE representation for $(p_{k}^{\mathrm{pw},\theta},z_{k}^{\theta},\tilde{z}_{k}^{\theta})$ whose drift matches the Euler discretization of the $\theta$-conditional Pontryagin adjoint BSDE associated with (53)–(54). The same argument applies to the derivatives $(p_{x,k}^{\mathrm{pw},\theta},p_{y,k}^{\mathrm{pw},\theta})$: they satisfy linearized discrete backward recursions obtained by differentiating the discrete adjoint equations, hence admit analogous discrete-BSDE representations with coefficients uniformly controlled on $K$.
@@
Step 4: Passage to continuous time and identification with the PMP costate. For each fixed $\theta\in K$, the forward SDE and the $\theta$-conditional adjoint BSDE form a standard FBSDE with coefficients parametrized by $\theta$. Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the continuous-time $\theta$-conditional Pontryagin objects under policy $\pi_{\varphi}$, so $p_{T}^{\theta}=U^{\prime}(X_{T}^{\pi,\theta})$.
Define the piecewise-constant interpolations
$p_{t}^{\Delta t,\theta}:=p_{k}^{\mathrm{pw},\theta},\qquad p_{x,t}^{\Delta t,\theta}:=p_{x,k}^{\mathrm{pw},\theta},\qquad p_{y,t}^{\Delta t,\theta}:=p_{y,k}^{\mathrm{pw},\theta},\hskip 18.49988ptt\in[t_{k},t_{k+1}).$
@@
By the same stability and convergence arguments as in huh2025breaking (Euler convergence for the forward equation plus discrete-BSDE convergence for the backward equation), we obtain, for each fixed $\theta\in K$,
@@
as $\Delta t\to 0$. Because all Lipschitz, growth, ellipticity, and covariance constants were assumed uniform on $K$, the convergence constants can be chosen independently of $\theta\in K$. This yields the claimed BPTT–PMP correspondence conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$, completing the proof. ∎
Appendix C Auxiliary results for Theorem 3
C.1 Stability of the projection map $(A,G)\mapsto-A^{-1}G$

### v2

@@ Appendix A Online uncertainty updates: Kalman–Bucy filtering and a plug-in decision-time benchmark
Remarks (relation to belief-aware control). The plug-in rule (85) is intentionally decision-time: it conditions on the current uncertainty description and does not attempt to optimize over how the posterior will evolve. In the present paper, we therefore treat the uncertainty law as fixed at a decision time (either as a fixed $q$ over a horizon, or as an externally updated sequence of inputs $q_{t}$ that is not controlled by the agent), mirroring the decision-time perspective common in Bayesian/learning portfolio-choice studies (Barberis, 2000; Pástor, 2000; Xia, 2001). Even in linear–Gaussian regimes where the belief state is finite-dimensional, the fully optimal partial-observation portfolio problem would treat the belief state (here, $(\hat{Y}_{t},P_{t})$) as part of the controlled state and optimize the policy in that enlarged state space (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017). Related necessary conditions under partial information can also be expressed via partial-observation maximum principles (Haussmann, 1987; Li and Tang, 1995; Baghery and Øksendal, 2007). Developing a belief-aware Pontryagin-guided policy optimizer that operates directly in $(x,y,\hat{Y},P)$-space (or its sufficient-statistic analogues) is an important direction that we defer to future work.
Appendix B Why dynamic programming and deep PDE surrogates are not primary baselines under the $\theta$-blind fixed-$q$ interface
This appendix explains why we do not treat classical dynamic programming (DP/HJB) or value-function-based deep PDE surrogates (PINNs / deep BSDE methods) as practical baselines in the high-dimensional uncertain markets targeted in this paper. DP is conceptually sound in low-dimensional Markovian settings (Fleming and Soner, 2006; Pham, 2009), but two obstructions dominate in our regime: (i) value-function approaches require accurate derivatives (often including mixed sensitivities) to recover optimal portfolios, and this becomes prohibitive as dimension and nonlinearity grow; (ii) principled parameter uncertainty magnifies both the state dimension and the aggregation difficulties. These points are standard for grid-based HJB solvers (Bellman, 1961; Kushner and Dupuis, 2001) and remain relevant for modern deep PDE surrogates (Raissi et al., 2019; Sirignano and Spiliopoulos, 2018; Han et al., 2018; Beck et al., 2019), especially under partial information/belief-state formulations (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
B.1 The value-function target for latent $\theta$ is a belief-state functional
Fixed-$q$ ex–ante objective and $\theta$-blind deployability.
Our decision-time interface is
$\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}J(\pi)\;:=\;\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}\mathbb{E}_{\theta\sim q}\!\left[\mathbb{E}\!\left[U(X_{T}^{\pi})\mid\theta\right]\right],$ (86)
where $\theta\sim q$ is latent, fixed along each episode, and $\mathcal{A}^{\mathrm{blind}}$ denotes admissible controls measurable with respect to the observable filtration (so the deployed policy cannot take $\theta$ as an input).
Why $(t,X_{t},Y_{t})$ is not a closed DP state.
In general, the observable history carries information about $\theta$, so the continuation value at time $t$ depends on the conditional law
$q_{t}(\cdot):=\mathcal{L}\!\big(\theta\mid\mathcal{F}_{t}^{\mathrm{obs}}\big),$ (87)
not only on $(t,X_{t},Y_{t})$. This is the standard partially observed control phenomenon: DP becomes Markov only after augmenting the state by a belief (or a finite-dimensional sufficient statistic when available) (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
Belief-state DP is infinite-dimensional under our interface.
The principled DP value function is therefore a functional of a probability measure:
$V(t,x,y,\mu):=\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}\mathbb{E}^{t,x,y,\mu}\!\left[U(X_{T}^{\pi})\right],\hskip 18.49988pt\mu\in\mathcal{P}(\Theta),$ (88)
so the associated HJB equation is posed on a space of measures in general (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017). Under our inference-agnostic interface, $q$ need not belong to a conjugate family and we do not specify a filtering model that would keep $q_{t}$ in a tractable finite-dimensional class. Consequently, a DP/PDE baseline would be forced to approximate $V$ over an infinite-dimensional belief state, which is not the computational target of this paper.
B.2 Even with deterministic parameters, HJB solvers face a derivative bottleneck in high dimension
Classical HJB: curse of dimensionality and full nonlinearity.
With deterministic parameters, DP leads to an HJB for $V(t,x,y)$ (Fleming and Soner, 2006; Pham, 2009). Grid-based solvers scale exponentially in the state dimension (Bellman, 1961; Kushner and Dupuis, 2001). In portfolio problems with $d$ assets and $m$ factors, the natural state already has dimension $m+2$, so even modest discretizations require $N^{m+2}$ grid points.
Moreover, realistic features (constraints, transaction costs, non-affine dynamics) typically yield fully nonlinear HJBs. Recovering the optimal portfolio from a value function also requires derivatives (schematically, in Merton-type problems)
$\pi^{\star}(t,x,y)\;\propto\;-\frac{V_{x}(t,x,y)}{V_{xx}(t,x,y)}\times(\text{risk model})^{-1}\times(\text{drift model}),$ (89)
and in factor-driven settings intertemporal hedging demands are driven by mixed sensitivities (e.g. $V_{xy}$). Thus the numerical burden is not only approximating $V$ but also obtaining stable first/second/mixed derivatives in high dimension.
B.3 Deep PDE surrogates remove grids but not the value-function derivative bottleneck
PINNs and deep BSDE methods replace grids with neural approximators trained on sampled points/paths (Raissi et al., 2019; Sirignano and Spiliopoulos, 2018; Han et al., 2018; Beck et al., 2019), but for fully nonlinear portfolio HJBs they remain value-function-based: they must implicitly learn high-dimensional gradients/Hessians and, crucially, mixed sensitivities that drive intertemporal hedging. In practice this can induce nonconvex and ill-conditioned training objectives (due to control suprema and nonlinear derivative dependence) and training signals that do not reliably control the specific derivative components needed for stable high-dimensional portfolio policies.
This is not a claim that deep PDE methods are “invalid”; rather, under our high-dimensional and $\theta$-blind interface, they are not practical baselines relative to direct, simulation-only policy optimization methods.
B.4 Why $\theta$-conditional value functions do not solve the $\theta$-blind ex–ante problem
A natural alternative is to condition on $\theta$ and solve a family of full-information problems.
$\theta$-conditional HJB.
If $\theta$ were observed, the full-information value function $V^{\theta}$ solves
$0=\partial_{t}V^{\theta}(t,x,y)+\sup_{\pi\in\mathbb{R}^{d}}\Big\{\mathcal{L}^{\pi,\theta}V^{\theta}(t,x,y)\Big\},\hskip 18.49988ptV^{\theta}(T,x,y)=U(x),$ (90)
for the controlled generator $\mathcal{L}^{\pi,\theta}$ (Fleming and Soner, 2006; Pham, 2009). This yields a $\theta$-dependent optimizer $\pi^{\star}(t,x,y,\theta)$, which is infeasible in our deployable $\theta$-blind setting.
Supremum–expectation non-commutativity.
More fundamentally, even perfect knowledge of the $\theta$-conditional optimizer does not solve the $\theta$-blind objective, because
$\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}\mathbb{E}_{\theta\sim q}\!\left[J(\pi,\theta)\right]\;\leq\;\mathbb{E}_{\theta\sim q}\!\left[\sup_{\pi}J(\pi,\theta)\right],\hskip 18.49988ptJ(\pi,\theta):=\mathbb{E}\!\left[U(X_{T}^{\pi})\mid\theta\right],$ (91)
and the inequality is typically strict. The right-hand side corresponds to the (infeasible) $\theta$-adaptive controller.
B.5 Why “sample $\theta$ only in the PDE loss” is mismatched to the ex–ante objective
One might try to train a single value network $V_{\varphi}(t,x,y)$ that does not take $\theta$ as input, and sample $\theta\sim q$ only inside a collocation/pathwise residual loss:
$\min_{\varphi}\;\mathbb{E}_{\theta\sim q}\!\Big[\big\|\mathrm{HJB}_{\theta}\!\big(V_{\varphi}\big)(t,x,y)\big\|^{2}\Big],\hskip 18.49988pt\text{(schematic)}$ (92)
where $\mathrm{HJB}_{\theta}(\cdot)$ is the full-information HJB operator for parameter $\theta$. This is appealing but generally mismatched to (86):
  • Different $\theta$ impose different (often conflicting) optimality conditions at the same $(t,x,y)$, so there is generally no single $V$ satisfying $\mathrm{HJB}_{\theta}(V)=0$ for many $\theta$ simultaneously.
  • Minimizing an expected squared PDE residual (92) is not equivalent to maximizing expected terminal utility under $\theta$-blind controls.
  • Portfolio feedback rules are typically extracted through nonlinear operations on derivatives (ratios, inverses), so small value-function approximation errors can induce large control errors.
B.6 A toy calculation: $\mathbb{E}[\pi^{\star}(\theta)]\neq\pi_{q}^{\star}$ under CRRA
The aggregation issue is visible even in a one-asset Merton model with static latent drift. Let
$\frac{dS_{t}}{S_{t}}=r\,dt+\theta\,dt+\sigma\,dW_{t},\hskip 18.49988pt\theta\sim\mathcal{N}(m,p),$
and restrict to constant fractions $\pi$ over a remaining horizon $\tau$. For CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$ with $\gamma>1$, the $\theta$-conditional full-information constant Merton rule is
$\pi^{\star}(\theta)=\frac{\theta}{\gamma\sigma^{2}}.$ (93)
Averaging this infeasible rule gives $\mathbb{E}[\pi^{\star}(\theta)]=m/(\gamma\sigma^{2})$. By contrast, the deployable $\theta$-blind ex–ante optimizer within the constant-fraction class is
$\pi_{q}^{\mathrm{const}}(\tau)=\frac{m}{\gamma\sigma^{2}+(\gamma-1)\tau\,p},\hskip 18.49988pt(\gamma>1),$ (94)
which exhibits shrinkage due to estimation uncertainty ($p>0$) and depends on the horizon $\tau$. Thus $\mathbb{E}[\pi^{\star}(\theta)]\neq\pi_{q}^{\mathrm{const}}(\tau)$ whenever $p>0$.
B.7 Implication for baselines in this paper
The discussion above motivates our baseline choices in the main text:
  • We validate on controlled drift/premium benchmarks where closed-form decision-time Gaussian targets are available and match the fixed-$q$ deployability restriction.
  • For learning-based comparisons, we emphasize simulation-only direct policy search baselines (e.g., PPO) that operate under the same observation restriction (no access to $\theta$; $\theta$ sampled only inside the simulator), rather than value-function PDE/BSDE solvers whose natural DP target is belief-augmented.
Appendix C Proof of Theorem 2
@@
Theorem 2 extends the BPTT–PMP (equivalently, BPTT–BSDE) correspondence established for deterministic-parameter models in our prior work on PG–DPO (see the main BPTT–BSDE correspondence result and proof in Huh et al. (2025b)). Here the only substantive change is that the market coefficients are indexed by a random but frozen parameter $\theta\sim q$, and we need convergence statements that hold conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$.  (This appendix replaces the current Appendix C in the draft.)
@@
Important remark (what this proof does not use). This proof concerns the $\theta$-conditional Pontryagin adjoint/costate for the fixed-$\theta$ control problem induced by (45)–(46). It does not use the $\theta$-blind $q$-aggregated stationarity condition (Theorem 1 in Section 3.2). Those constructions affect only stage 2 projection targets and are irrelevant to the BPTT–PMP convergence itself.
@@
Notation, filtration, and the pathwise-vs-adapted issue. Fix $\theta\in\Theta$. We work conditionally on this $\theta$ and consider the augmented (simulator) filtration
@@
$\mathbb{G}^{\theta}:=(G_{t}^{\theta})_{t\in[0,T]},\hskip 18.49988ptG_{t}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t\}\big)\ \text{(with the usual augmentation)}.$
All conditional expectations and $L^{2}$ projections below are taken with respect to $G_{t_{k}}^{\theta}$.
Key point. The raw BPTT “adjoints” are pathwise derivatives of the terminal utility along the full discrete computation graph. Hence, for $k<N$, they are typically $G_{T}^{\theta}$-measurable and need not be $G_{t_{k}}^{\theta}$-measurable (they can depend on future increments). In contrast, the stochastic maximum principle (SMP/PMP) costate is $\mathbb{G}^{\theta}$-adapted. Therefore, the correspondence is formulated after a standard one-step conditional $L^{2}$ projection (equivalently, a discrete-time BSDE representation), which turns pathwise objects into adapted discrete adjoints.
@@
Let $\Delta t>0$, $t_{k}:=k\Delta t$, $k=0,\dots,N$, $N\Delta t=T$. For readability we suppress the policy parameters $\varphi$ and write $\pi_{k}:=\pi_{\varphi}(t_{k},X_{k}^{\theta},Y_{k}^{\theta})$, where $\pi_{\varphi}$ is $\theta$-blind but evaluated along the $\theta$-conditional trajectory.
@@
Step 1: Conditioning on $\theta$ and uniformity of bounds. Fix a compact set $K\subset\Theta$. Assume the coefficients in (45)–(46) satisfy the usual Lipschitz and linear-growth conditions uniformly over $\theta\in K$, and that the block covariance structure of $(W,W^{Y})$ (including instantaneous correlation) is uniformly nondegenerate on $K$. Assume also the regularity/integrability needed for the $\theta$-conditional SMP (including the blocks relevant for the portfolio Hamiltonian gradient) to be well posed. Then, for each fixed $\theta\in K$, the controlled SDE system is well posed and admits uniform-in-time $L^{2}$ moment bounds. Moreover, the Euler–Maruyama scheme enjoys the standard strong error bound
$\sup_{t\in[0,T]}\mathbb{E}\!\left[\left\|(X_{t}^{\pi,\theta},Y_{t}^{\theta})-(X_{t}^{\Delta t,\theta},Y_{t}^{\Delta t,\theta})\right\|^{2}\right]^{1/2}\leq C_{K}\,\Delta t^{1/2},$
@@
with a constant $C_{K}$ that can be chosen independently of $\theta\in K$. These are exactly the deterministic assumptions used in Huh et al. (2025b), now stated uniformly on $K$.
@@
Step 2: Discrete forward scheme and BPTT pathwise adjoints (fixed $\theta$). Under fixed $\theta$, consider the Euler scheme on the grid $(t_{k})$:
@@
with terminal reward $U(X_{N}^{\theta})$. Define the discrete pathwise wealth costate
@@
and the additional pathwise blocks
$p_{x,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial X_{k}^{\theta}},\hskip 18.49988ptp_{y,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial Y_{k}^{\theta}}.$
Automatic differentiation/BPTT computes $\{(p_{k}^{\mathrm{pw},\theta},p_{x,k}^{\mathrm{pw},\theta},p_{y,k}^{\mathrm{pw},\theta})\}_{k=0}^{N}$ via the backward chain rule along the discrete forward graph. The one-step algebraic form of the backward recursion coincides with the deterministic-parameter analysis in Huh et al. (2025b), with the replacements
@@
and with the factor block handled exactly as in the wealth–factor extension therein. All one-step remainder terms are controlled by standard Taylor/Euler estimates with constants uniform in $\theta\in K$.
Step 3: Conditional $L^{2}$ projection $\Rightarrow$ adapted discrete BSDE variables (fixed $\theta$). Fix $\theta\in K$. Let $\Delta W_{k}:=W_{t_{k+1}}-W_{t_{k}}$ and $\Delta W_{k}^{Y}:=W^{Y}_{t_{k+1}}-W^{Y}_{t_{k}}$. Uniform nondegeneracy of the block covariance of $(\Delta W_{k},\Delta W_{k}^{Y})$ implies that the conditional $L^{2}$-projection onto $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ is well defined (and yields unique coefficients).
(a) One-step projection (defining the discrete martingale coefficients). Define $(z_{k}^{\theta},\tilde{z}_{k}^{\theta})$ by the conditional orthogonality relations (equivalently, the $L^{2}$-projection) such that
@@
$p_{k+1}^{\mathrm{pw},\theta}=\mathbb{E}\!\left[p_{k+1}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right]+z_{k}^{\theta}\Delta W_{k}+\tilde{z}_{k}^{\theta}\Delta W_{k}^{Y}+\varepsilon_{k+1}^{\theta},$
where $\varepsilon_{k+1}^{\theta}$ is orthogonal (in conditional $L^{2}$) to $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ given $G_{t_{k}}^{\theta}$.
(b) Adapted discrete adjoints (the objects compared to PMP). Define the adapted discrete-time costates by conditional expectation:
$p_{k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{k}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right],\hskip 18.49988ptp_{x,k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{x,k}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right],\hskip 18.49988ptp_{y,k}^{\Delta t,\theta}:=\mathbb{E}\!\left[p_{y,k}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right].$
These are $G_{t_{k}}^{\theta}$-measurable and are the discrete-time counterparts of the $\theta$-conditional Pontryagin objects.
@@
(c) Discrete BSDE form and identification of the driver. Substituting the projection in (a) into the one-step BPTT recursion from Step 2 and then taking conditional expectations with respect to $G_{t_{k}}^{\theta}$ yields a canonical discrete-time BSDE representation for the adapted triple $(p_{k}^{\Delta t,\theta},z_{k}^{\theta},\tilde{z}_{k}^{\theta})$. The resulting drift coincides with the Euler discretization of the $\theta$-conditional adjoint BSDE associated with (45)–(46) under the fixed policy $\pi_{\varphi}$, up to standard one-step remainder terms of higher order in $\Delta t$. The same argument applies to the blocks $(p_{x,k}^{\Delta t,\theta},p_{y,k}^{\Delta t,\theta})$: they satisfy linearized discrete backward recursions (obtained by differentiating the discrete adjoint equations) and hence admit analogous discrete-BSDE representations with coefficients uniformly controlled on $K$. All estimates and remainder controls are identical to Huh et al. (2025b) once we condition on $\theta$, and the constants can be chosen uniformly on $K$ by the uniform-in-$\theta$ assumptions from Step 1.
@@
Step 4: Passage to continuous time and convergence to the $\theta$-conditional PMP costate. For each fixed $\theta\in K$, the forward SDE and the $\theta$-conditional adjoint BSDE form a standard FBSDE with coefficients parametrized by $\theta$. Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the continuous-time $\theta$-conditional Pontryagin objects under $\pi_{\varphi}$, with terminal condition $p_{T}^{\theta}=U^{\prime}(X_{T}^{\pi,\theta})$.
Define the piecewise-constant interpolations of the adapted discrete variables:
$p_{t}^{\Delta t,\theta}:=p_{k}^{\Delta t,\theta},\qquad p_{x,t}^{\Delta t,\theta}:=p_{x,k}^{\Delta t,\theta},\qquad p_{y,t}^{\Delta t,\theta}:=p_{y,k}^{\Delta t,\theta},\hskip 18.49988ptt\in[t_{k},t_{k+1}).$
@@
Then, by the same stability and convergence arguments as in Huh et al. (2025b) (Euler convergence for the forward equation plus convergence of the discrete-BSDE scheme for the backward equation), we obtain, for each fixed $\theta\in K$,
@@
as $\Delta t\to 0$. Because all Lipschitz, growth, ellipticity, and block-covariance constants were assumed uniform on $K$, the constants in these convergence bounds can be chosen independently of $\theta\in K$. This yields the claimed BPTT–PMP correspondence conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$, completing the proof. ∎
Appendix D Auxiliary results for Theorem 3
D.1 Stability of the projection map $(A,G)\mapsto-A^{-1}G$